# Bytes to Schlep? Use a FEP: Hiding Protocol Metadata with Fully Encrypted Protocols

### v1

@@ 5.2. Properties
The channel construction in Figure 2 satisfies correctness for the secure close function $\mathscr{C}\equiv 0$, Traffic Shaping, FEP-CCFA, INT-CST, and IND-CCFA.
@@ 6.4.1. Confidentiality and Integrity definitions
We adopt confidentiality and integrity definitions from Bellare and Namprempre (Bellare and Namprempre, 2008). We require only slight modifications to adapt them to our stateful communication channel, and the details appear in Appendix B.
@@ 6.4.2. Fully Encrypted Datagram Protocols
A channel satisfies FEP-$x$, for $x\in\{\textrm{CPA},\textrm{CCA}\}$ if, for a security parameter $\lambda$ and PPT adversary $\mathcal{A}$, $\left|P\left[\mathsf{Exp}^{\textsf{FEP-$x$},b}_{\mathcal{A}}(1^{\lambda})=1|b\overset{{\scriptscriptstyle\operatorname{R}}}{\leftarrow}\{0,1\}\right]-1/2\right|$ is negligible in $\lambda$.
@@ 8.1. The Construction
In our construction, we assume the same AEAD scheme from Section 2 and, as before, drop its associated data argument. Let $\ell_{\textrm{Overhead}}=\ell_{\textrm{Nonce}}+\ell_{\textrm{Tag}}$ represent the encryption overhead since our construction uses fresh nonces for each ciphertext. We give an explicit encoding scheme for our message space into bytes as follows: we encode the null message $\top$ as a single zero byte, and prefix all other messages (including the empty string $\epsilon$) with a leading $1$ byte. We set $\ell_{\textrm{out}}=65507$ (the maximum size UDP payloand length), and we include in our message space all byte strings of length up to $\ell_{\textrm{in}}=\ell_{\textrm{out}}-\ell_{\textrm{Overhead}}-3$, reserving two bytes for the plaintext length and one for the message type. We denote the length of a ciphertext containing the null message as $\ell_{\textrm{null}}=1+\ell_{\textrm{Overhead}}$. We abuse notation and write $|m|$ to mean both the size of $m$ and the two byte unsigned integer that represents that size, depending on the context.
@@ 8.2. Properties
The channel construction in Figure 3 satisfies correctness for Datagram channels, Traffic Shaping, FEP-CCA, INT-CTXT, and IND-CCA.
@@ 9. Existing Fully Encrypted Protocols
Obfs4/Lyrebird Auth Fail X Padding $44^{*}$
@@
Since our work is motivated the lack of security definitions for existing Fully Encrypted Protocols, we identified a set of protocols that make an attempt to appear fully encrypted above the transport layer for analysis. Our goal in analyzing these protocols is to determine if these protocols satisfy our definitions, and to what extent they have identifying features that our security definitions are designed to address. We excluded closed-source protocols but note that many existing implementations of FEPs that do not appear in our list nonetheless adapt the approach of one of the protocols we analyze (outlinevpn, 2024; Psiphon, 2024; Tunnelblick, 2023). We did not perform a thorough audit of the security engineering of the protocols’ implementations, which we consider outside the scope of our paper.
@@ 9.2. Results
We first describe the eight FEP implementations studied: (1) Shadowsocks (sha, 2022) is a fully encrypted SOCKS5 proxy with many implementations and the ability to proxy both datastream and datagram traffic. We examined the Shadowsocks-libev (Windy et al., 2020) implementation in both TCP and UDP configurations. (2) We also studied a Shadowsocks implementation in the censorship-circumvention software suite V2Ray (v2ray, 2024). (3) VMess (vmess, 2024) is a custom FEP also implemented inside V2Ray for censorship circumvention. (4) Obfs4 (obfs4, 2024) is a FEP designed for Tor (Dingledine et al., 2004). (5) OpenVPN (OpenVPN, 2024) is open-source VPN software that functions in datastream or datagram mode, and there is a well-known patch available designed to obfuscate its traffic called the XOR patch (OpenVPN XORPatch, 2018). (6) Obfuscated SSH is modification of OpenSSH that fully encrypts its handshake messages, and it can be configured with a Pre-Shared Key. (7) kcptun is a FEP designed to be reliable and fast over very noisy networks. (8) Wireguard-SWGP (swg, 2024) is a proxy for Wireguard (Donenfeld, 2017) that fully encrypts its traffic.
@@
We also experimentally observed that the protocols in the V2Ray framework fail to satisfy datastream integrity (and thus cannot satisfy FEP-CCFA), dropping isolated messages silently when certain ciphertext bytes were modified in transit. We reported this as a security vulnerability to the project maintainers on January 4, 2024 and the issue was resolved in version 5.14.1.
@@ 10. Discussion
Unfortunately, as we have shown in Section 9, existing FEPs and their implementations can be effectively distinguished in practice because of a failure to satisfy one or more of the security definitions we present in this work. We consider our definitions a major step towards articulating and achieving this goal for FEPs, but using them to guarantee FEP indistinguishability requires some further determinations be made that we consider outside the scope of our paper. In particular, a secure close function must be selected and a Traffic Shaping schedule or distribution must be determined. Further, some variability in these selections may be necessary for practical functionality and efficiency reasons.
@@ 11. Related Work
We build on the model of Fischlin et al. (Fischlin et al., 2015) to study data streams in the presence of fragmentation, a problem also been studied elsewhere (Boldyreva et al., 2012; Albrecht et al., 2009, 2019, 2016). Fragmentation is a particular for FEPs, as plaintext length fields cannot be used. The notion of Boundary Hiding (Boldyreva et al., 2012) is related to but does not imply our FEP notions because, for one, it does not enforce random-appearing outputs.
Some aspects of Fully Encrypted Protocols have been studied, such as active probing (Frolov et al., 2020; Beznazwy and Houmansadr, 2020), and attacks on confidentiality (Ji et al., 2022; Peng, 2020). Much research has focused on detecting FEPs (Wang et al., 2015; Wu et al., 2023; Frolov and Wustrow, 2019; Xue et al., 2022) using a variety of approaches. Fifield (Fifield, 2023) outlined a series of implementation weaknesses in existing FEPs. Bernstein et al. (Bernstein et al., 2013) developed a method to encode elliptic curve points as uniformly random strings.
@@
Some FEPs designs have come from elsewhere than the open-source community. Obfs4/Lyrebird is based on Scramblesuit (Winter et al., 2013). InterMAC, which is a near-FEP, has been both analyzed (Boldyreva et al., 2012) and implemented (Albrecht et al., 2019). The IETF proposal for a Pseudorandom Extension of cTLS (Schwartz and Patton, 2022) suggest a FEP encoding of TLS to improve protocol security and privacy rather than for censorship circumvention.
@@ 12. Conclusion and Future Work
A natural enhancement to our constructions would be a fully encrypted key exchange for forward secrecy. Obfs4 (obfs4, 2024) uses Elligator (Bernstein et al., 2013) for this purpose, but it remains to prove its security as well as investigate other techniques. Relatedly, one could formulate a notion of forward metadata secrecy, where the FEP properties may still be preserved even if a long-term key is compromised. In addition, FEP notions should be explored in models where securely creating shared state is not trivial but bidirectional communication is possible.
@@ Acknowledgments
This work was supported by the Office of Naval Research (ONR) and the Defense Advanced Research Projects Agency (DARPA). We would also like to thank Ben Robinson, who helped perform the Wireguard experiments, and the anonymous referees who reviewed this paper for their helpful suggestions and advice.
@@ Appendix E Datastream Construction Proofs
We follow bytes as they pass through the three ciphertext buffers: first, every call of Send either returns immediately or appends a length-payload block pair to obuf. By the length regularity of the AEAD scheme, $\ell_{c}$ is the size of the ciphertext constructed in Line 20, and the correctness of the padding clear by inspection. Then, every fragment $c$ removed from obuf is removed from the front of the buffer, and is returned by Send, meaning that $C\|(c\|\texttt{obuf})=(C\|c)\|\texttt{obuf}$ remains fixed. Similarly, bytes are removed from $C$ and passed into $\texttt{buf}_{R}$ by the same mechanism, fixing the concatenated ciphertext buffers. Finally, we show that Recv only removes full length-payload block pairs from Recv. But this is clear from the conditions in Lines 7, 12, 15: Recv waits to receive a full length block, and removes it only if its following payload block is empty, or the full payload block has been received (in which case the payload block is removed as well). ∎
@@
Next consider a call $\textsc{Send}(m,p,f)$. First, $m$ is concatenated to $\texttt{buf}_{S}$, which preserves our invariant since it is removed from the beginning of $M_{s}$ and placed in the end of $\texttt{buf}_{S}$, leaving $\texttt{buf}_{S}\|M_{s}$ fixed. If Send returns via the conditional in Line 3, bytes are transferred from obuf to $C$, leaving $C\|\texttt{obuf}$ fixed. We consider two cases, where $\ell_{c}=0$ and otherwise and in each case calculate the new value of $I$, labeled $I^{\prime}$. If $\ell_{c}=0$, $\texttt{buf}_{S}$ is unchanged, $c$ is a single block, produced in Line 15, and we note that $r(c)=\epsilon$. Thus, in this case after Send returns, we have a new value for our invariant, $I^{\prime}=M_{r}\|r(\texttt{buf}_{R}\|C\|\texttt{obuf}\|c)\|\texttt{buf}_{S}\|M_{s}$, which equal to the previous $I$. If $\ell_{c}>0$, $c$ consists of two ciphertext blocks, and $r(c)=\epsilon\|\texttt{buf}_{S}[1..o]$. Then we can write:
@@
Since our construction is structurally similar to the AEAD construction in Fischlin et al. (Fischlin et al., 2015), we adopt the same strategy for our proof and we reduce the INT-CST game to the AUTH game for our underlying AEAD encryption scheme by assuming there is some adversary $A$ which wins the INT-CST game against our construction with advantage $\epsilon$, and constructing an adversary $B$ that runs $A$ and simulates the behavior of Send and Recv, replacing the encryption and decryption calls with their corresponding oracles in the AUTH experiment. For the details, we refer to Fischlin et al. (Fischlin et al., 2015), and we note where our proof differs. We note that the sequence of blocks recorded by $B$ are entirely ciphertext blocks, in contrast to the previous approach which interleaved plaintext length fields between each ciphertext block, however the sequence numbers are structured identically, incrementing by one with each block. Thus, $B$ records ciphertext blocks when they are passed to Send and “decrypts” each block when a genuine ciphertext fragment is passed to Recv. The argument is the same up until the first non-genuine ciphertext fragment $c^{\prime}$ is passed to Recv. $B$ extracts the genuine ciphertext blocks from the beginning of $c^{\prime}$ and ”decrypts” them. We note that for our channel, blocks can produce empty plaintext upon decryption and indeed that an AUTH forgery can take place within a ciphertext block that causes Recv to produce nothing (e.g., if a length block is forged, or a payload block full of padding), but that violating INT-CST requires that $A$ produce inputs to Recv that cause it to generate a nonempty plaintext fragment. This must, by construction of Recv, occur within a fragment of ciphertext which is parsed by Recv as a payload block with non zero length, i.e. in Line 18. This means that the plaintext fragment to be produced by Recv as a result of deviating ciphertext, $m^{\prime}$ must be the result of decrypting a payload block, and so if a deviation occurs within the preceeding length block, this is already a forgery and $B$ can wait to collect $\ell_{\texttt{len}}$ bytes to produce the forged length block; by assumption, if Recv is to produce a ciphertext, Dec will produce some valid nonnegative integer when applied to the block or Recv sets $\texttt{fail}=1$, precluding future outputs. If the first deviation occurs within a payload block, the preceeding length block is genuine and has length $\ell$ and so $B$ can produce $\epsilon$ until $\ell$ bytes have been received, producing the full ciphertext block forgery. Thus, $B$ can construct the forgery for our protocol and INT-CST is satisfied.

### v2

@@ 5.2. Properties
If the AEAD scheme (Gen, Enc, Dec) in the construction in Figure 2 satisfies IND\$-CPA, INT-CTXT, and is length additive (see Sec 2) then the channel construction satisfies correctness for the secure close function $\mathscr{C}\equiv 0$, Traffic Shaping, FEP-CCFA, INT-CST, and IND-CCFA.
@@ 6.4.1. Confidentiality and Integrity definitions
We adopt confidentiality and integrity definitions from Bellare and Namprempre (Bellare and Namprempre, 2008). We require only slight modifications to adapt them to our stateful communication channel. Details appear in Appendix B.
@@ 6.4.2. Fully Encrypted Datagram Protocols
A channel satisfies FEP-$x$, for $x\in\{\textrm{CPA},\textrm{CCA}\}$ if, for a security parameter $\lambda$ and PPT adversary $\mathcal{A}$, $\left|P\left[\mathsf{Exp}^{\textsf{FEP-$x$},b}_{\mathcal{A}}(1^{\lambda})=1\big{|}b\overset{{\scriptscriptstyle\operatorname{R}}}{\leftarrow}\{0,1\}\right]-1/2\right|$ is negligible in $\lambda$.
@@ 8.1. The Construction
In our construction, we assume the same AEAD scheme from Section 2. For simplicity, we further assume that Enc includes in its outputs each nonce as a prefix of the ciphertext when it is called, and that Dec extracts the first $\ell_{\textrm{Nonce}}$ bytes from its input as the nonce to use for decryption. Let $\ell_{\textrm{Overhead}}=\ell_{\textrm{Nonce}}+\ell_{\textrm{Tag}}$ represent the encryption overhead since our construction uses fresh nonces for each ciphertext. We give an explicit encoding scheme for our message space into bytes as follows: we encode the null message $\top$ as a single zero byte, and prefix all other messages (including the empty string $\epsilon$) with a leading $1$ byte. We set $\ell_{\textrm{out}}=65507$ (the maximum size UDP payload length), and we include in our message space all byte strings of length up to $\ell_{\textrm{in}}=\ell_{\textrm{out}}-\ell_{\textrm{Overhead}}-3$, reserving two bytes for the plaintext length and one for the message type. We denote the length of a ciphertext containing the null message as $\ell_{\textrm{null}}=1+\ell_{\textrm{Overhead}}$. We abuse notation and write $|m|$ to mean both the size of $m$ and the two byte unsigned integer that represents that size, depending on the context.
Our protocol is stateless, so each function simply returns the symmetric key $k$ as the new state after each call. The Send function generates a fresh nonce, and if Traffic Shaping is off, directly applies Enc to produce the datagram. If the desired output length $p$ is too small for an authenticated message, we produce random bytes (Lines 6–7). We produce arbitrary length chaff messages (Line 9) or padded plaintext messages (Line 13), and return with an error if $p$ is too large to fit in a datagram or $m$ is too large to fit within $p\geq 0$ (Line 10). Recv interprets all small messages as chaff, and then decrypts larger messages, returning the plaintext, chaff symbol, or error where appropriate.
@@ 8.2. Properties
If the AEAD scheme (Gen, Enc, Dec) in the channel construction in Figure 3 satisfies IND\$-CPA, INT-CTXT, and length additivity (see Sec 2) then the channel construction satisfies correctness for Datagram channels, Traffic Shaping, FEP-CCA, INT-CTXT, and IND-CCA.
@@ 9. Existing Fully Encrypted Protocols
Obfs4 Auth Fail X Padding $44^{*}$
@@
Since our work is motivated the lack of security definitions for existing Fully Encrypted Protocols, we identified a set of protocols that make an attempt to appear fully encrypted above the transport layer for analysis. Our goal in analyzing these protocols is to determine if these protocols satisfy our definitions, and to what extent they have identifying features that our security definitions are designed to address. We excluded closed-source protocols but note that many existing implementations of FEPs that do not appear in our list nonetheless adapt the approach of one of the protocols we analyze (Outline Client., 2024; Psiphon., 2024; Tunnelblick., 2023). We did not perform a thorough audit of the security engineering of the protocols’ implementations, which we consider outside the scope of our paper.
@@ 9.2. Results
We first describe the eight FEP implementations studied: (1) Shadowsocks (Shadowsocks., 2022) is a fully encrypted SOCKS5 proxy with many implementations and the ability to proxy both datastream and datagram traffic. We examined the Shadowsocks-libev (Windy et al., 2020) implementation in both TCP and UDP configurations. (2) We also studied a Shadowsocks implementation in the censorship-circumvention software suite V2Ray (Project V., 2024). (3) VMess (VMess., 2024) is a custom FEP also implemented inside V2Ray for censorship circumvention. (4) Obfs4 (obfs4, The obfourscator) is a FEP designed for Tor (Dingledine et al., 2004). (5) OpenVPN (OpenVPN Project., 2024) is open-source VPN software that functions in datastream or datagram mode, and there is a well-known patch available designed to obfuscate its traffic called the XOR patch (OpenVPN XORPatch., 2018). (6) Obfuscated SSH is modification of OpenSSH that fully encrypts its handshake messages, and it can be configured with a Pre-Shared Key. (7) kcptun is a FEP designed to be reliable and fast over very noisy networks. (8) Wireguard-SWGP (Simple Wireguard Proxy., 2024) is a proxy for Wireguard (Donenfeld, 2017) that fully encrypts its traffic.
@@
We also experimentally observed that the protocols in the V2Ray framework fail to satisfy datastream integrity (and thus cannot satisfy FEP-CCFA), dropping isolated messages silently when certain ciphertext bytes were modified in transit. We reported this as a security vulnerability to the project maintainers on January 4, 2024 and the issue was resolved in version 5.14.1.222The bug was caused by a failure to propogate errors properly when decryption fails during the data transport phase.
@@ 10. Discussion
Modeling limitations
Our modeling choices limit our results in some ways. We follow existing models for stateful communication channels, which, for simplicity, provide unidirectional rather than bidirectional communication and omit time. Bidirectional communication can be achieved within our model by composing two unidirectional channels, one in each direction. However, such channels are not allowed to share state, including keys and other initialization parameters, and not observing this restriction can violate the security guarantees. Such a limitation reduces potential efficiency over the truly bidirectional setting. Similarly, omitting time from our model precludes some convenient protocol features, such as timeouts leading to channel closures and time-based replay protections. This limits both the design of new protocols and the analysis of existing ones. However, security models have been designed that include time (Backes et al., 2014; Canetti et al., 2017), and our models and constructions can be extended to work within such a setting.
FEP indistinguishability
@@
Unfortunately, as we have shown in Section 9, existing FEPs and their implementations can be effectively distinguished in practice because of a failure to satisfy one or more of the security definitions we present in this work. We consider our definitions a major step towards articulating and achieving this goal for FEPs, but using them to guarantee FEP indistinguishability requires some further determinations to be made that we consider outside the scope of our paper. In particular, a secure close function must be selected and a Traffic Shaping schedule or distribution must be determined. Further, some variability in these selections may be necessary among FEPs for practical functionality and efficiency reasons.
Secure close functions
@@
Recommendations for FEP designers
We intend our results to inform the design of existing and future FEPs. Many of the existing FEPs could satisfy our novel security definitions with straightforward changes to the message formats and protocol implementation. We make the following specific recommendations regarding the two novel protocol features we introduce, secure close functions and Traffic Shaping:
  1. (1)
    We suggest ideally adopting the close function of InterMAC, where the receiver closes after error only at a multiple of $n$ bytes, for constant $n$, by adding a MAC at those positions. These MACs would be in addition to any MACs needed to support variable-length records and could occur at relatively low frequency, where such frequency should be chosen to be acceptable across applications and protocols to maximize FEP indistinguishability (e.g. every 10K bytes). We further suggest closing the connection after a certain amount of time has passed with no ciphertext received, which is technically outside our model. The combination of these close behaviors would permit connections to eventually close after an error is observed without adding too much bandwidth overhead. An alternative suggestion is simply to adopt a connection timeout after not receiving any ciphertext. This behavior does permit an active adversary to keep a connection open by continually sending traffic, but it is a more minor change to implement.
  2. (2)
    To implement Traffic Shaping, datastream FEPs should be extended to buffer content that does not fit within the prescribed output length $p$ and to produce padding when $p$ exceeds the amount necessary for the message. Datagram FEPs should be extended to include the use of null messages and to refuse to send messages that cannot fit within the prescribed length. FEP developers should thus implement Traffic Shaping, although existing applications need not necessarily use it and instead await more work on appropriate traffic patterns.
@@ 11. Related Work
We build on the model of Fischlin et al. (Fischlin et al., 2015) to study data streams in the presence of fragmentation, a problem also been studied elsewhere (Boldyreva et al., 2012; Albrecht et al., 2009, 2019, 2016). Fragmentation is a problem for FEPs, as plaintext length fields cannot be used. The notion of Boundary Hiding (Boldyreva et al., 2012) is related to but does not imply our FEP notions because, for one, it does not enforce random-appearing outputs.
Some aspects of Fully Encrypted Protocols have been studied, such as active probing (Frolov et al., 2020; Alice et al., 2020), and attacks on confidentiality (Ji et al., 2022; Peng, 2020). Much research has focused on detecting FEPs (Wang et al., 2015; Wu et al., 2023; Frolov and Wustrow, 2019; Xue et al., 2022) using a variety of approaches. Fifield (Fifield, 2023) outlined a series of implementation weaknesses in existing FEPs. Bernstein et al. (Bernstein et al., 2013) developed a method to encode elliptic curve points as uniformly random strings.
@@
Some FEPs designs have come from elsewhere than the open-source community. Obfs4 (obfs4, The obfourscator) is based on Scramblesuit (Winter et al., 2013) and has been forked under the new name Lyrebird (Lyrebird., 2024). InterMAC, which is a near-FEP, has been both analyzed (Boldyreva et al., 2012) and implemented (Albrecht et al., 2019). The IETF proposal for a Pseudorandom Extension of cTLS (Schwartz and Patton, 2022) suggest a FEP encoding of TLS to improve protocol security and privacy rather than for censorship circumvention.
@@ 12. Conclusion and Future Work
A natural enhancement to our constructions would be a fully encrypted key exchange for forward secrecy. Obfs4 (obfs4, The obfourscator) uses Elligator (Bernstein et al., 2013) for this purpose, but it remains to prove its security as well as investigate other techniques. Relatedly, one could formulate a notion of forward metadata secrecy, where the FEP properties may still be preserved even if a long-term key is compromised. In addition, FEP notions should be explored in models where securely creating shared state is not trivial but bidirectional communication is possible.
@@ Acknowledgments
This work was supported by the Office of Naval Research (ONR) and the Defense Advanced Research Projects Agency (DARPA). We would also like to thank Ben Robinson, who helped perform the Wireguard experiments, and David Fifield and the anonymous referees who reviewed this paper for their helpful suggestions and advice.
@@ Appendix E Datastream Construction Proofs
We follow bytes as they pass through the three ciphertext buffers: first, every call of Send either returns immediately or appends a length-payload block pair to obuf. By the length regularity of the AEAD scheme, $\ell_{c}$ is the size of the ciphertext constructed in Line $17$, and the correctness of the padding clear by inspection. Then, every fragment $c$ removed from obuf is removed from the front of the buffer, and is returned by Send, meaning that $C\|(c\|\texttt{obuf})=(C\|c)\|\texttt{obuf}$ remains fixed. Similarly, bytes are removed from $C$ and passed into $\texttt{buf}_{R}$ by the same mechanism, fixing the concatenated ciphertext buffers. Finally, we show that Recv only removes full length-payload block pairs from Recv. But this is clear from the conditions in Lines 6, 7, 11: Recv waits to receive a full length block, and removes only when the full following payload block has been received (in which case the payload block is removed as well). ∎
@@
Next consider a call $\textsc{Send}(m,p,f)$. First, $m$ is concatenated to $\texttt{buf}_{S}$, which preserves our invariant since it is removed from the beginning of $M_{s}$ and placed in the end of $\texttt{buf}_{S}$, leaving $\texttt{buf}_{S}\|M_{s}$ fixed. If Send returns via the conditional in Line 3, bytes are transferred from obuf to $C$, leaving $C\|\texttt{obuf}$ fixed. Otherwise, a new fragment $c$ is produced and added to obuf. We calculate the new value of $I$, labeled $I^{\prime}$. $c$ consists of two ciphertext blocks, and $r(c)=\epsilon\|\texttt{buf}_{S}[1..o]$. Then we can write:
@@
Since our construction is structurally similar to the AEAD construction in Fischlin et al. (Fischlin et al., 2015), we adopt the same strategy for our proof and we reduce the INT-CST game to the AUTH game for our underlying AEAD encryption scheme by assuming there is some adversary $A$ which wins the INT-CST game against our construction with advantage $\epsilon$, and constructing an adversary $B$ that runs $A$ and simulates the behavior of Send and Recv, replacing the encryption and decryption calls with their corresponding oracles in the AUTH experiment. For the details, we refer to Fischlin et al. (Fischlin et al., 2015), and we note where our proof differs. We note that the sequence of blocks recorded by $B$ are entirely ciphertext blocks, in contrast to the previous approach which interleaved plaintext length fields between each ciphertext block, however the sequence numbers are structured identically, incrementing by one with each block. Thus, $B$ records ciphertext blocks when they are passed to Send and “decrypts” each block when a genuine ciphertext fragment is passed to Recv. The argument is the same up until the first non-genuine ciphertext fragment $c^{\prime}$ is passed to Recv. $B$ extracts the genuine ciphertext blocks from the beginning of $c^{\prime}$ and ”decrypts” them. We note that for our channel, blocks can produce empty plaintext upon decryption and indeed that an AUTH forgery can take place within a ciphertext block that causes Recv to produce nothing (e.g., if a length block is forged, or a payload block full of padding), but that violating INT-CST requires that $A$ produce inputs to Recv that cause it to generate a nonempty plaintext fragment. This must, by construction of Recv, occur within a fragment of ciphertext which is parsed by Recv as a payload block with non zero length, i.e. in Line 14. This means that the plaintext fragment to be produced by Recv as a result of deviating ciphertext, $m^{\prime}$ must be the result of decrypting a payload block, and so if a deviation occurs within the preceeding length block, this is already a forgery and $B$ can wait to collect $\ell_{\texttt{len}}$ bytes to produce the forged length block; by assumption, if Recv is to produce a ciphertext, Dec will produce some valid nonnegative integer when applied to the block or Recv sets $\texttt{fail}=1$, precluding future outputs. If the first deviation occurs within a payload block, the preceeding length block is genuine and has length $\ell$ and so $B$ can produce $\epsilon$ until $\ell$ bytes have been received, producing the full ciphertext block forgery. Thus, $B$ can construct the forgery for our protocol and INT-CST is satisfied.